1. On Sep 26th 2016 you sold a put option of $3(put premium) to give the buyer the right to sell a share of facebook on Oct 9th, 2016 at the strike price, X=$120. At the same time , you also sold short a facebook share, So is $128. If facebook shares go up to $140 on Oct 9th, what is your profit or loss ?

2. On sep 26th 2016, you sold a put option of $4 (premium) to give the buyer the right to sell a share of facebook on Nov 18th, 2016 at the strike price , X=$123. The current facebook share, So is $128 If facebook shares go up to $135 on Nov 18th, what is your profit or loss ?
Answer #1

1.
Profit=-140+128-MAX(120-140,0)+3=-9

Loss of $9

2.
Profit=-135+128-MAX(123-135,0)+3=-4
Loss of $4